The cost of new machine will be $100000 after 12 years. In that time the old machine can be sold for $5000 as salvage value. What amount should be retained out of profits at the end of each year accumulate at C.I at 5% P.A?
1
Expert's answer
2020-09-22T16:23:56-0400

To get the amount to be retained out of profits per year

="\dfrac{cost . of asset- salvageValue}{time \space useful}"

= "\dfrac{100000- 50000}{12}"

="\dfrac{95000}{12}"

=$ 7916.67